x=3^1/2,y=5^1/3 then (x+y)’s rationalise factor

Grade:

1 Answers

Y RAJYALAKSHMI
9 years ago
First, make the powers equal
x = 31/2 = 33/6 = 271/6
y = 51/2 = 52/6 = 251/6
x + y  = 271/6 + 251/6
Now, rationalising factor is  (271/6 - 251/6) (272/3 + 271/3251/3252/3)